OpenStudy (anonymous):

Model each problem with a quadratic equation. Then solve. 17. Find the side of a square with an area of 625 ft^2. 18.Find the radius of a circle with an area of 153.86 ft^2.

OpenStudy (anonymous):

@onemillion6

OpenStudy (anonymous):

17. If a square has side s, what is the formula for it's area? 18. Similar to 17, what is the formula for the radius of a circle with radius r? Once you have the formulas, you can set them equal to what the question gives you and solve!

OpenStudy (anonymous):

Not just going to give you the answer :\

OpenStudy (haseeb96):

@adoredelano look here the area of square formula is length multiply with width ( l x w ) so 

OpenStudy (anonymous):

25

OpenStudy (anonymous):

Area = x^2 = 625 x^2 = 625 x^2-625 = 0 (x-25)(x+25)= 0 Positive answer: x = 25

OpenStudy (haseeb96):